# The Transition from Giant Planets to Brown Dwarfs beyond 1 au from the Stellar Metallicity Distribution

### v1

@@ Abstract
journal: the Astronomical Journalsoftware: PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), scikit-learn (Pedregosa et al., 2011)
@@ 1 Introduction
In the past, the classification of an object as a giant planet or a brown dwarf was solely dependent on its mass in relation to the deuterium-burning limit ($\sim 13\,M_{\rm Jup}$; Spiegel et al. 2011; Bodenheimer et al. 2013): those with masses below the limit were considered giant planets and those with masses above the limit were considered brown dwarfs (Boss et al., 2007). More recently, it has been argued that the classifications of substellar companions should instead depend on formation mechanism, where only objects that form “like planets” are considered planets (e.g., Chabrier et al., 2014). Objects that form like planets are those that form bottom-up, beginning with the aggregation of a solid core onto which gas is accreted from the surrounding protoplanetary disk. This core-accretion scenario has long stood as the most likely formation mechanism for gas giant planets and is believed to be able to form objects with masses several times that of Jupiter (Pollack et al., 1996; Alibert et al., 2005; Hubickyj et al., 2005; Matsuo et al., 2007). Conversely, objects that form top-down, through the direct collapse of gas without a preexisting solid core, are considered to form “like stars” (Bate, 2012; Kratter & Lodato, 2016; Offner et al., 2023). Despite this definition, top-down formation mechanisms are thought to be capable of forming companions with masses comparable to that of Jupiter in certain environments (e.g., Diamond & Parker, 2024; Fu et al., 2025). The top-down formation mechanism of disk fragmentation (also referred to as “gravitational instability”), for instance, can form substellar companions with Jupiter-like and greater masses at wide orbital separations (Adams & Benz, 1992; Boss, 1997; Durisen et al., 2007; Kratter et al., 2010; Zhu et al., 2012). These predictions make it challenging to determine how substellar companions form based on their masses alone.
@@
A number of studies have explored the differences in the metallicity distributions of giant planet and brown dwarf host stars using a variety of methods and observational samples. Ma & Ge (2014) used a Kolmogorov-Smirnov test to show that companions with masses between $1\,M_{\rm Jup}$ and $5\,M_{\rm Jup}$ orbit stars with a distinct metallicity distribution from those with masses above $13\,M_{\rm Jup}$. Evidence for the presence of two distinct host star metallicity distributions were also found by Mata Sánchez et al. (2014) and Maldonado & Villaver (2017). Santos et al. (2017) used a multivariate Gaussian clustering technique on a sample of companions primarily detected by radial velocity surveys and found evidence for a transition in the stellar metallicity distribution around a minimum mass ($M{\rm sin}(i)$, where $M$ is the mass of the companion and $i$ is the inclination of the orbit) of $4\,M_{\rm Jup}$, where less massive companions tend or orbit stars that are more metal rich. This analysis was performed using homogeneously determined stellar metallicities from the SWEET-Cat database (Santos et al., 2013) and companions with orbital periods mostly under 1 yr (i.e., orbital periods under 365 days for $88.75\%$ of the sample). Schlaufman (2018, hereafter S18) performed a similar analysis using transiting giant planets, brown dwarfs, and low-mass stars with close-in orbits (i.e., semi-major axes $a<0.25$ au for $95\%$ of the sample) and found that the transition between populations most likely occurs between $4$ and $10\,M_{\rm Jup}$. Most recently, Wang (2025) investigated the metallicities of wide-separation ($a>10$ au), directly imaged giant planets and brown dwarfs and found that companions with masses under $10\,M_{\rm Jup}$ tend to have super-stellar metallicities and companions with masses over $20\,M_{\rm Jup}$ tend to have metallicies similar to their host stars, suggesting a transition between $10$ and $20\,M_{\rm Jup}$.
@@
In the present paper, we revisit the question of whether formation mechanisms for giant planets and brown dwarfs can be distinguished via the companion mass – stellar metallicity distribution. Our analysis differs from previous studies in a few ways. First, we use a sample of uniformly characterized systems from a single, robustly assembled observational survey. This approach allows us to avoid biases introduced by compiling a sample of objects from different surveys, which may have employed different target selection functions and analysis techniques. Second, our sample consists only of objects with a=150a=1-50 au, probing a distinct range of orbital separations compared to previous works. Lastly, we utilize a hierarchical Bayesian approach for modeling the transition between two distinct stellar metallicity distributions as a function of companion mass. Previous techniques, by comparison, relied on transiting or directly imaged planets to estimate the transition mass and have only been capable of estimating a singular value or a simple range of values. For complementary analyses of the transition between giant planets and brown dwarfs based on orbital eccentricities and occurrence rates, we direct the reader to G Gilbert et al. (2025) and J. Van Zandt et al. (in review).
@@ 2 Sample
Recently, J. Van Zandt et al. (submitted) generated posterior distributions for the true masses of 192 of the companions detected in CLS (all companions with $M_{c}{\rm sin}(i)<80\,M_{\rm Jup}$) by performing joint orbital fits on the radial velocity data and Gaia/Hipparcos astrometry. We used these $M_{c}$ posterior distributions in our analysis, such that we can avoid needing to marginalize over orbital inclination.
We selected all systems that have companions with $0.5\leq M_{c}/M_{\rm Jup}\leq 110$ and with $a\geq 1$ au. Our imposed lower limit on $M_{c}$ was selected to avoid including companions that were unlikely to have formed via core accretion. In addition, by limiting our sample to very massive objects, we avoid potential biases caused by survey incompleteness; the CLS dataset is nearly $100\%$ sensitive to companions in this mass range (see Rosenthal et al. 2021 and J. Van Zandt et al. submitted). Our imposed limits on $a$ were chosen because the CLS found very few objects with $M_{c}>10\,M_{\rm Jup}$ within 1 au. The limits therefore restricted the sample to orbital separations in which companions spanning nearly the full mass range were detected. Lastly, following J. Van Zandt et al. (submitted), we excluded HD 167215 c from the analysis, due to the companion being a likely false positive. In addition, we omit HIP 57050 d and HD 45184 d due to these companions having highly unconstrained masses. Our sample of 85 companions is shown in Figure 1. We note that all of the stars in the sample have $T_{\rm eff}<6500$ K and $80\%$ having $T_{\rm eff}$ between 5000 and 6200 K. This is an artifact of the selection function used by CLS, which avoided hot stars due to rapid rotations that limit radial velocity precision.
@@ 3.1 Hierarchical Bayesian Model
also referred to as a “change-point” model (Raftery & Akman, 1986; Green, 1995). Here, $\mathcal{N}$ indicates a Normal distribution and $M_{c}$ is sampled directly from the posterior distribution generated by J. Van Zandt et al. (submitted). The observed metallicity ${\rm[Fe/H]}_{o}^{j}$ is modeled as a random Gaussian variable with the form
@@
We sampled our model via Markov-Chain Monte Carlo (MCMC) using PyMC (Abril-Pla et al., 2023). We ran the MCMC with 10 chains for $120,000$ steps each, discarding the first $20,000$ steps as burn-in. We ensured the convergence of the model using the Gelman-Rubin statistic ($\hat{r}<1.1$; Gelman & Rubin 1992).
@@ 3.3 Tests for Statistical Significance
3.4 Evidence of Distinct Eccentricity Distributions
Formation mechanism is also believed to influence orbital eccentricity; companions that form bottom-up are predicted to have relatively low orbital eccentricities compared to companions that form top-down (e.g., Veras et al., 2009; Bate, 2012; Duffell & Chiang, 2015; Grishin & Perets, 2015). We tested this prediction by comparing the eccentricity distributions on either end of our best-fit value of $\gamma$, which are shown in Figure 5. It was visually evident that the low-mass population has a higher probability of being in low-eccentricity orbits, whereas the high-mass population has a mean orbital eccentricity of $\sim 0.5$. A two-sample Anderson-Darling test of the two distributions (divided again at $M_{c}$ of $27\,M_{\rm Jup}$) yielded a test statistic value of 4.05, allowing us to reject the null hypothesis that the two samples come from the same distribution to a significance level $99\%$. A two-sample Kolmogorov-Smirnov test yielded a p-value of $0.03$, providing similar evidence. Thus, we found that the orbital eccentricity distributions of the two populations are indeed statistically distinct. Nonetheless, we emphasize that just because the two eccentricity distributions are distinct when divided at 27 $M_{\rm Jup}$, it does not mean that this is the mass where the eccentricity distributions change. G. Gilbert et al. (2025) performed a more careful analysis of the eccentricities in the CLS sample and find that the eccentricity distribution changes gradually, rather than sharply, with increasing companion mass. We refer the reader to G. Gilbert et al. (2025) for more information on that analysis.
3.5 Application to S18 Sample
As a comparison, we ran a modified version of our HBM on the sample from S18. The primary change to the model was the removal of the inclination term and the replacement of $M{\rm sin}(i)$ with mass, which was possible because all of the companions in S18 are transiting. This analysis yielded a transition mass of $\gamma=59^{+10}_{-25}\,M_{\rm Jup}$, significantly higher than the value of $\sim 10\,M_{\rm Jup}$ reported by S18. For the [Fe/H] distributions, we found $C_{1}=0.12^{+0.02}_{-0.01}$, $S_{1}=0.10\pm 0.01$, $C_{2}=-0.12^{+0.06}_{-0.07}$, and $S_{2}=0.16^{+0.04}_{-0.03}$. These results are displayed in Figure 3.
@@ 3.6 Clustering Analysis
As an independent check of our results, we repeated the three clustering analyses in S18. Specifically, we utilized a hierarchical clustering algorithm (scipy.cluster.hierarchy.linkage), a k-means clustering algorithm (scipy.cluster.vq.kmeans), and Gaussian mixture clustering algorithm (sklearn.mixture.GaussianMixture). We performed each calculation $10^{3}$ times, each time resampling $M_{c}$ and [Fe/H] $10^{3}$ according to the observed values and their uncertainties. With each iteration, we recorded the most massive companion classified as belonging to the low-mass population, which we treated as a proxy for the transition mass.
After all $10^{3}$ realizations were complete, we calculated the medians and $1\sigma$ uncertainties of the resulting transition mass distributions. The hierarchical clustering algorithm returned a relatively uninformative transition mass distribution of $6^{+14}_{-2}\,M_{\rm Jup}$. The k-means clustering algorithm predicted a relatively low transition masses of $6^{+4}_{-1}\,M_{\rm Jup}$. Lastly, the Gaussian mixture clustering algorithm predicted a transition mass of and $20^{+5}_{-4}\,M_{\rm Jup}$. These values span a wide range and are systematically lower than the values returned by out HBM analysis, as is the case for the S18 sample. We believe that these discrepancies are due to an over-sensitivity of the clustering algorithms to the locations of individual data points near the transition region. The HBM is less biased to the locations of individual data points in $M_{c}-{\rm[Fe/H]}$ space, instead focusing on where the overall [Fe/H] distribution changes.
@@ 4.1 Occurrence Rates
The earliest estimates of $\gamma$ were made based on the occurrence rates of giant planets and brown dwarfs. Grether & Lineweaver (2006) analyzed the $M{\rm sin}(i)$ distribution of substellar companions detected by radial velocity surveys and found that it reaches a minimum at $M{\rm sin}(i)=31^{+25}_{-18}\,M_{\rm Jup}$. Sahlmann et al. (2011) later broke the orbital inclination degeneracy for this sample using Hipparcos astrometry and found that the minimum occurs between true masses of $25\,M_{\rm Jup}$ and $45\,M_{\rm Jup}$. This finding agrees with our estimate of $\gamma$, suggesting that the transition in formation mechanism supported by our analysis of stellar metallicities may coincide with a local minimum in occurrence rate. For an analysis of the occurrence rates of giant planets and brown dwarfs in the CLS sample, see J. Van Zandt et al. (in review).
@@ 4.2 Orbital Architectures
The orbits of substellar companions also provide useful clues towards their formations. In particular, bottom-up formation is expected to produce relatively circular orbits compared to top-down formation due to interactions between the protoplanetary disk and the embedded companion (e.g., Veras et al., 2009; Bate, 2012; Duffell & Chiang, 2015; Grishin & Perets, 2015). In addition, bottom-up formation is predicted to result in orbits aligned with the stellar equator more frequently than top-down formation (e.g., Bate et al., 2010; Offner et al., 2016). A number of studies have investigated these predictions for eccentricity (Ma & Ge, 2014; Bowler et al., 2020; Vowell et al., 2025) and spin-orbit angle (Bryan et al., 2020, 2021; Rice et al., 2022; Bowler et al., 2023; Giacalone et al., 2024; Poon et al., 2024; Rusznak et al., 2025; Radzom et al., 2024; Smith et al., 2024; Wang et al., 2024, e.g.,). Estimates for $\gamma$ based on orbital eccentricity range from $10\,M_{\rm Jup}$ to several tens of $M_{\rm Jup}$. Our somewhat simple analysis of orbital eccentricities finds support for a transition in eccentricities around ${\sim}27\,M_{\rm Jup}$ (Figure 5), while a more careful analysis of the CLS eccentricity distribution in G. Gilbert et al. (2025) finds evidence for a gradual transition. We were refer the reader to that paper for a more in-depth discussion of this topic. Analyses of spin-orbit angles have found that close-in companions are more likely to be aligned with the equators of their host stars when higher in mass (Giacalone et al., 2024; Radzom et al., 2024), whereas analyses of wide-separation companions have found the opposite (Bowler et al., 2023). Evidently, the question of how to distinguish bottom-up from top-down formation on the basis of orbital architectures remains far from settled.

### v2

@@ Abstract
software: PyMC (Abril-Pla et al., 2023), SciPy (Virtanen et al., 2020), scikit-learn (Pedregosa et al., 2011)
@@ 1 Introduction
In the past, the classification of an object as a giant planet or a brown dwarf was solely dependent on its mass in relation to the deuterium-burning limit ($\sim 13\,M_{\rm Jup}$; Saumon et al. 1996; Chabrier et al. 2000; Spiegel et al. 2011; Bodenheimer et al. 2013): those with masses below the limit were considered giant planets and those with masses above the limit were considered brown dwarfs (Boss et al., 2007). However, it has been argued that the classifications of substellar companions should instead depend on formation mechanism, where only objects that form “like planets” are considered planets (e.g., Burrows et al., 2001; Luhman, 2008; Chabrier et al., 2014). Objects that form like planets are those that form bottom-up, beginning with the aggregation of a solid core onto which gas is accreted from the surrounding protoplanetary disk. This core-accretion scenario has long stood as the most likely formation mechanism for gas giant planets and is believed to be able to form objects with masses several times that of Jupiter (Pollack et al., 1996; Alibert et al., 2005; Hubickyj et al., 2005; Matsuo et al., 2007). Conversely, objects that form top-down, through the direct collapse of gas without a preexisting solid core, are considered to form “like stars” (Bate, 2012; Kratter & Lodato, 2016; Offner et al., 2023). Despite this definition, top-down formation mechanisms are thought to be capable of forming companions with masses comparable to that of Jupiter in certain environments (e.g., Diamond & Parker, 2024; Fu et al., 2025). The top-down formation mechanism of disk fragmentation (also referred to as “gravitational instability”), for instance, can form substellar companions with Jupiter-like and greater masses at wide orbital separations (Adams & Benz, 1992; Boss, 1997; Durisen et al., 2007; Kratter et al., 2010; Zhu et al., 2012). These predictions make it challenging to determine how substellar companions form based on their masses alone.
@@
A number of studies have explored the differences in the metallicity distributions of giant planet and brown dwarf host stars using a variety of methods and observational samples. Ma & Ge (2014) used a Kolmogorov-Smirnov test to show that companions with masses between $1\,M_{\rm Jup}$ and $5\,M_{\rm Jup}$ orbit stars with a distinct metallicity distribution from those with masses above $13\,M_{\rm Jup}$. Evidence for the presence of two distinct host star metallicity distributions were also found by Mata Sánchez et al. (2014) and Maldonado & Villaver (2017). Santos et al. (2017) used a multivariate Gaussian clustering technique on a sample of companions primarily detected by radial velocity surveys and found evidence for a transition in the stellar metallicity distribution around a minimum mass ($M_{c}{\rm sin}(i)$, where $M_{c}$ is the mass of the companion and $i$ is the inclination of the orbit) of $4\,M_{\rm Jup}$, where less massive companions tend or orbit stars that are more metal rich. This analysis was performed using homogeneously determined stellar metallicities from the SWEET-Cat database (Santos et al., 2013) and companions with orbital periods mostly under 1 yr (i.e., orbital periods under 365 days for $88.75\%$ of the sample). Schlaufman (2018, hereafter S18) performed a similar analysis using transiting giant planets, brown dwarfs, and low-mass stars with close-in orbits (i.e., semi-major axes $a<0.25$ au for $95\%$ of the sample) and found that the transition between populations most likely occurs between $4$ and $10\,M_{\rm Jup}$. Most recently, Wang (2025) investigated the metallicities of wide-separation ($a>10$ au), directly imaged giant planets and brown dwarfs and found that companions with masses under $10\,M_{\rm Jup}$ tend to have super-stellar metallicities and companions with masses over $20\,M_{\rm Jup}$ tend to have metallicies similar to their host stars, suggesting a transition between $10$ and $20\,M_{\rm Jup}$.
@@
In the present paper, we revisit the question of whether formation mechanisms for giant planets and brown dwarfs can be distinguished via the companion mass – stellar metallicity distribution. Our analysis differs from previous studies in a few ways. First, we use a sample of uniformly characterized systems from a single, robustly assembled observational survey. This approach allows us to avoid biases introduced by compiling a sample of objects from different surveys, which may have employed different target selection functions and analysis techniques. Second, our sample consists only of objects with a=150a=1-50 au, probing a distinct range of orbital separations compared to previous works. Lastly, we utilize a hierarchical Bayesian approach for modeling the transition between two distinct stellar metallicity distributions as a function of companion mass. Previous techniques, by comparison, relied on transiting or directly imaged planets to estimate the transition mass and have only been capable of estimating a singular value or a simple range of values. For complementary analyses of the transition between giant planets and brown dwarfs based on orbital eccentricities and occurrence rates, we direct the reader to Gilbert et al. (2025) and Van Zandt et al. (2025).
@@ 2 Sample
Recently, Van Zandt et al. (2025) generated posterior distributions for the true masses of 192 of the companions detected in CLS (all companions with $M_{c}{\rm sin}(i)<80\,M_{\rm Jup}$) by performing joint orbital fits on the radial velocity data and Gaia/Hipparcos astrometry. We used these $M_{c}$ posterior distributions in our analysis, such that we can avoid needing to marginalize over orbital inclination.
We selected all systems that have companions with $0.5\leq M_{c}/M_{\rm Jup}\leq 110$ and with $a\geq 1$ au. Our imposed lower limit on $M_{c}$ was selected to avoid including companions that were unlikely to have formed via core accretion. In addition, by limiting our sample to very massive objects, we avoid potential biases caused by survey incompleteness; the CLS dataset is nearly $100\%$ sensitive to companions in this mass range (see Rosenthal et al. 2021 and Van Zandt et al. 2025). Our imposed limits on $a$ were chosen because the CLS found very few objects with $M_{c}>10\,M_{\rm Jup}$ within 1 au. The limits therefore restricted the sample to orbital separations in which companions spanning nearly the full mass range were detected. Lastly, following Van Zandt et al. (2025), we excluded HD 167215 c from the analysis, due to the companion being a likely false positive. In addition, we omit HIP 57050 d and HD 45184 d due to these companions having highly unconstrained masses. Our sample of 85 companions is shown in Figure 1. We note that all of the stars in the sample have $T_{\rm eff}<6500$ K and $80\%$ having $T_{\rm eff}$ between 5000 and 6200 K. This is an artifact of the selection function used by CLS, which avoided hot stars due to rapid rotations that limit radial velocity precision.
@@ 3.1 Hierarchical Bayesian Model
also referred to as a “change-point” model (Raftery & Akman, 1986; Green, 1995). Here, $\mathcal{N}$ indicates a Normal distribution and $M_{c}$ is sampled directly from the posterior distribution generated by Van Zandt et al. (2025). The observed metallicity ${\rm[Fe/H]}_{o}^{j}$ is modeled as a random Gaussian variable with the form
@@
We sampled our model via Markov-Chain Monte Carlo (MCMC) using PyMC (Abril-Pla et al., 2023). We ran the MCMC with 10 chains for $120,000$ steps each, discarding the first $20,000$ steps as burn-in. We ensured the convergence of the model using the Gelman-Rubin statistic ($\hat{r}<1.1$; Gelman & Rubin 1992.
@@ 3.3 Tests for Statistical Significance
3.4 Evidence of Distinct Eccentricity Distributions
Formation mechanism is also believed to influence orbital eccentricity; companions that form bottom-up are predicted to have relatively low orbital eccentricities compared to companions that form top-down (e.g., Veras et al., 2009; Bate, 2012; Duffell & Chiang, 2015; Grishin & Perets, 2015). We tested this prediction by comparing the eccentricity distributions on either end of our best-fit value of $\gamma$, which are shown in Figure 5. It was visually evident that the low-mass population has a higher probability of being in low-eccentricity orbits, whereas the high-mass population has a mean orbital eccentricity of $\sim 0.5$. A two-sample Anderson-Darling test of the two distributions (divided again at $M_{c}$ of $27\,M_{\rm Jup}$) yielded a test statistic value of 4.05, allowing us to reject the null hypothesis that the two samples come from the same distribution to a significance level $99\%$. A two-sample Kolmogorov-Smirnov test yielded a p-value of $0.03$, providing similar evidence. Thus, we found that the orbital eccentricity distributions of the two populations are indeed statistically distinct. Nonetheless, we emphasize that just because the two eccentricity distributions are distinct when divided at 27 $M_{\rm Jup}$, it does not mean that this is the mass where the eccentricity distributions change. Gilbert et al. (2025) performed a more careful analysis of the eccentricities in the CLS sample and find that the eccentricity distribution changes gradually, rather than sharply, with increasing companion mass. We refer the reader to Gilbert et al. (2025) for more information on that analysis.
3.5 Application to S18 Sample
As a comparison, we ran a modified version of our HBM on the sample from S18. The primary change to the model was the removal of the inclination term and the replacement of $M_{c}{\rm sin}(i)$ with mass, which was possible because all of the companions in S18 are transiting. This analysis yielded a transition mass of $\gamma=59^{+10}_{-25}\,M_{\rm Jup}$, significantly higher than the value of $\sim 10\,M_{\rm Jup}$ reported by S18. For the [Fe/H] distributions, we found $C_{1}=0.12^{+0.02}_{-0.01}$, $S_{1}=0.10\pm 0.01$, $C_{2}=-0.12^{+0.06}_{-0.07}$, and $S_{2}=0.16^{+0.04}_{-0.03}$. These results are displayed in Figure 3.
@@ 3.6 Clustering Analysis
As an independent check of our results, we repeated the three clustering analyses in S18. Specifically, we utilized a hierarchical clustering algorithm (scipy.cluster.hierarchy.linkage), a k-means clustering algorithm (scipy.cluster.vq.kmeans), and Gaussian mixture clustering algorithm (sklearn.mixture.GaussianMixture). We performed each calculation $10^{3}$ times, each time resampling $M_{c}$ and [Fe/H] $10^{3}$ according to the observed values and their uncertainties. With each iteration, we recorded both the most massive companion assigned to the low-mass population and the least massive companion assigned to the high-mass population. To determine the effective $1\sigma$ confidence interval of the transition mass, we calculated the $16^{\rm th}$ percentile of the former distribution and the $84^{\rm th}$ percentile of the latter distribution.
The hierarchical clustering algorithm returned a transition mass range of 4 to 34 $M_{\rm Jup}$. The k-means clustering algorithm predicted a relatively low transition masses range of 2 to 11 $M_{\rm Jup}$. Lastly, the Gaussian mixture clustering algorithm predicted a transition mass range of 16 to 41 $M_{\rm Jup}$. These values span a wide range and generally agree with the transition mass predicted by our HBM, with the exception of the k-means approach.
@@ 4.1 Occurrence Rates
The earliest estimates of $\gamma$ were made based on the occurrence rates of giant planets and brown dwarfs. Grether & Lineweaver (2006) analyzed the $M_{c}{\rm sin}(i)$ distribution of substellar companions detected by radial velocity surveys and found that it reaches a minimum at $M_{c}{\rm sin}(i)=31^{+25}_{-18}\,M_{\rm Jup}$. Sahlmann et al. (2011) later broke the orbital inclination degeneracy for this sample using Hipparcos astrometry and found that the minimum occurs between true masses of $25\,M_{\rm Jup}$ and $45\,M_{\rm Jup}$. A similar minimum was found more recently by Currie et al. (2023). These findings agree with our estimate of $\gamma$, suggesting that the transition in formation mechanism supported by our analysis of stellar metallicities may coincide with a local minimum in occurrence rate. For an analysis of the occurrence rates of giant planets and brown dwarfs in the CLS sample, see Van Zandt et al. (2025).
@@ 4.2 Orbital Architectures
The orbits of substellar companions also provide useful clues towards their formations. In particular, bottom-up formation is expected to produce relatively circular orbits compared to top-down formation due to interactions between the protoplanetary disk and the embedded companion (e.g., Veras et al., 2009; Bate, 2012; Duffell & Chiang, 2015; Grishin & Perets, 2015). In addition, bottom-up formation is predicted to result in orbits aligned with the stellar equator more frequently than top-down formation (e.g., Bate et al., 2010; Offner et al., 2016). A number of studies have investigated these predictions for eccentricity (Ma & Ge, 2014; Bowler et al., 2020; Vowell et al., 2025a) and spin-orbit angle (e.g., Bryan et al., 2020, 2021; Rice et al., 2022; Bowler et al., 2023; Giacalone et al., 2024; Poon et al., 2024; Rusznak et al., 2025; Radzom et al., 2024; Smith et al., 2024; Wang et al., 2024). Estimates for $\gamma$ based on orbital eccentricity range from $10\,M_{\rm Jup}$ to several tens of $M_{\rm Jup}$. Our somewhat simple analysis of orbital eccentricities finds support for a transition in eccentricities around ${\sim}27\,M_{\rm Jup}$ (Figure 5), while a more careful analysis of the CLS eccentricity distribution in Gilbert et al. (2025) finds evidence for a gradual transition. We were refer the reader to that paper for a more in-depth discussion of this topic. Analyses of spin-orbit angles have found that close-in companions are more likely to be aligned with the equators of their host stars when higher in mass (Giacalone et al., 2024; Radzom et al., 2024; Carmichael et al., 2025; Vowell et al., 2025b), whereas analyses of wide-separation companions have found the opposite (Bowler et al., 2023). Evidently, the question of how to distinguish bottom-up from top-down formation on the basis of orbital architectures remains far from settled.